how far up the wall is a 13ft ladder if the base of the ladder is 8ft away from the wall?
Answer by nerdybill(7384)

how far up the wall is a 13ft ladder if the base of the ladder is 8ft away from the wall?
.
Applying Pythagorean theorem:
Let h= height up wall
then
h^2 + 8^2 = 13^2
h^2 + 64 = 169
h^2 = 105
h = 10.25 feet